OpenStudy (anonymous):

yes final project =sembreak ^^ obtain the positive solution less than 360 degree 1. sec^2 x-4=0 2. cot^2x=3 3. 2cos^2x=cosx 4. cot^2x=cotx 5. 4sin^2x-3=0

OpenStudy (anonymous):

congratz sembreak ^^ hope many will answer ^^ god bless ^^

OpenStudy (anonymous):

\[sec^2x - 4=0\]\[\Leftrightarrow\frac{1}{cos^2x}=4\]\[\Leftrightarrow \frac{1}{4}=cos^2x\]

OpenStudy (anonymous):

<> <<<<is this equivalent? yehey 4 more to go

OpenStudy (anonymous):

The \(\Leftrightarrow\) sign means that the two lines 'mutually imply' each other, which is a fancy way of saying that they effectively mean the same thing.

OpenStudy (anonymous):

Incidentally, what did you get for x when you solved that?

OpenStudy (anonymous):

from my project 1-40 this 5 question is 36-40.. so meaning i dont have an idea sory this is the hardest part

OpenStudy (anonymous):

Well if you have \[\frac{1}{4}=cos^2x\], then you can take the square root of both sides:\[\Leftrightarrow \pm\frac{1}{2}=cosx\] Which gives you 4 different solutions for x, each of which is equally valid. What are those 4 solutions?